Express the following as a complex number in the form a + ib:
The complex conjugate of i4n-3, with n Î Z
Please show you working out, and thank you in advance for your help

Found 2 solutions by AnlytcPhil, venugopalramana:
Answer by AnlytcPhil(1810)

Express the following as a complex number in the form a + ib:

The complex conjugate of i4n-3, with n Î Z

By definition we know that i2 = -1

We know that i4n = i2·2·n = ((i2)2)n = ((-1)2)n = (1)n = 1

Since i4 = 1

i4n-3 = i4n-3i4 = i4n-3+4 = i4n+1 = i4ni1 = 1·i = i

Now we want the conjugate of i

The conjugate of A + Bi is A - Bi and vice-versa

We write i as 0 + i, so its conjugate is 0 - i, or -i, so

Answer = -i

Edwin
AnlytcPhil@aol.com

Answer by venugopalramana(3286)

I^(4N-3)=I^4N/I^3={(I^2)^2N}/{(I^2)(I)}={(-1)^2}^N/{(-1)(I)=(1)^N/(-I)=-1/I
=(-1)(I)/{(I)(I)}=(-1)(I)/(-1)=I=0+1*I